OpenStudy (anonymous):

Which of the following is the slope between the two points (-5, -1)and (6, -4)?

OpenStudy (anonymous):

first correct answer= medal

OpenStudy (anonymous):

It's a repost.

OpenStudy (anonymous):

use slope equation m = y_2 - y_1 / x_2 - x_1 m = -1 - (-4) / -5 - 6 m = 3 / -11